A beach ball, moving at a speed of 1.27 m/s, rolls off a pier and hits the water 0.75 m away from the end of the pier. How high above the water is the pier?
1
Expert's answer
2012-10-15T11:54:19-0400
time of falling is
t = 0.75/1.27 = 0.59 s
height above the water is
h=gt^2/2 = 9.8*0.59^2/2 = 1.7 m